OpenStudy (anonymous):

In slope-intercept form, write the equation of the line that is parallel toy= -6x - 22 and passes through (-4,1)

OpenStudy (mathstudent55):

Parallel lines have the same slope. If you have the equation of a line in the form y = mx + b, then m is the slope. Line y = -6x - 22 is in y = mx + b form and m = slope = -6 Any line parallel to y = -6x - 22 has slope m = -6 The equation of a line through a point (x1, y1) and with slope m is given by \[y - y _{1} = m(x - x _{1})\] For the new line, you have the slope = m = -6, and you have (x1,y1) = (-4,1) Can you try it now?

OpenStudy (anonymous):

im confused??

OpenStudy (mathstudent55):

you have an equation when given slope and a point. y = y1 = m(x - x1) You have m and x1,y1. Insert the values of m, x1, and y1 into that equation, and all you'll heve will be x, y, and numbers. That's your equation.

OpenStudy (anonymous):

ohhh ok!! thank you

OpenStudy (anonymous):

i think i did it wrong........... ill show you my work y=y1=m(x-x1) y-1-=6(x-(-4) y=1=-6x-24 +6 +6 y=1=x-30 i did it wrong